# Joint Beamforming Optimization and Mode Selection for RDARS-aided MIMO Systems

### v1

@@ V Simulations
To determine the suitable value for the number of elements in the connection mode in terms of the hardware complexity and performance gain, we analyze the ANMSE performance versus the number of elements in the connection mode under various system typologies and channel conditions, as shown in Fig. 8. We consider four channel conditions based on the presence or absence of LoS components in the BS-UE and RDARS-UE channels. In topology 1, the RDARS is located in the middle of the BS and users, while in topology 2, it is closer to the BS. We clearly find that the performance gain of the proposed scheme is not significant as $a$ increases in both topologies when the RDARS-UE channel lacks LoS components. However, when the LoS paths exist between the RDARS and users, the RDARS architecture exhibits substantial performance improvement as $a$ increases. Furthermore, the slope of MSE performance initially increases and then decreases as the number of elements in the connection mode increases when the deterministic LoS paths exist in the RDARS-UE channel. This suggests an optimal number of connected elements for maximizing the slope of performance improvement exists. Specifically, when the BS-UE channel follows Rayleigh fading, the slope becomes larger when $a\leq 6$ and then decreases when $a>6$, where the transformation threshold is equal to the number of users. However, under the condition of the BS-UE channel subject to Rician fading, this slope changes at $a=2$, which equals the difference between the number of receive antennas at the BS and the number of users. In conclusion, the optimal number of elements in the connection mode on RDARS depends on the channel conditions, the number of users, and the number of receive antennas at the BS. These findings provide crucial insights when selecting the appropriate value of connected elements in the practical implementation of RDARS-aided systems.
@@
In Fig. 9, we explored the ANMSE performance versus the Rician factor $\kappa_{t}$ under different mode selection schemes with the same phase coefficients, considering both single-user and multiple-user scenarios. In the single-user case, as the Rician factor $\kappa_{t}$ increases from 0 to 1, the channel transitions from Rayleigh fading to a LoS channel. Consequently, the ANMSE performance is gradually improved. Additionally, as the Rician factor grows, the performance gap between the optimized mode selection scheme and the random/fixed mode selection schemes diminishes. This implies that the optimization of the mode selection becomes less critical as the strength of the LoS component increases. On the other hand, the ANMSE performance in the case of multiple users does not exhibit a monotonic relationship with the Rician factor due to the increased interference among multiple users as the LoS components become more dominant. The presence of stronger LoS components will lead to more significant interference, impacting the overall performance. Lastly, when considering only LoS components ($\kappa_{T}=1$), both the optimized mode selection scheme and the random/fixed mode selection scheme achieve the same ANMSE performance. This finding aligns with Proposition 3, as stated in Section III, suggesting the mode selection matrix ${\bf{A}}$ can be arbitrarily chosen when only LoS channels are present. In conclusion, the ANMSE performance in the RDARS is influenced by the Rician factor, with different trends observed in single-user and multiple-user scenarios. The necessity for mode selection optimization diminishes when only LoS components are present. These insights are valuable for guiding the practical implementation of RDARS systems.

### v2

@@ V Simulations
To determine the suitable value for the number of elements in the connection mode in terms of the hardware complexity and performance gain, we analyze the ANMSE performance versus the number of elements in the connection mode under various system typologies and channel conditions, as shown in Fig. 9. We consider four channel conditions based on the presence or absence of LoS components in the BS-UE and RDARS-UE channels. In topology 1, the RDARS is located in the middle of the BS and users with $d_{R}=100$, while in topology 2, it is closer to the BS with $d_{R}=200$. We clearly find that the performance gain of the proposed scheme is not significant as $a$ increases in both topologies when the RDARS-UE channel lacks LoS components. However, when the LoS paths exist between the RDARS and users, the RDARS architecture exhibits substantial performance improvement as $a$ increases. Furthermore, the slope of MSE performance initially increases and then decreases as the number of elements in the connection mode increases when the deterministic LoS paths exist in the RDARS-UE channel. This suggests an optimal number of connected elements for maximizing the slope of performance improvement exists. Specifically, when the BS-UE channel follows Rayleigh fading, the slope becomes larger when $a\leq 6$ and then decreases when $a>6$, where the transformation threshold is equal to the number of users. However, under the condition of the BS-UE channel subject to Rician fading, this slope changes at $a=2$, which equals the difference between the number of receive antennas at the BS and the number of users. In conclusion, the optimal number of elements in the connection mode on RDARS depends on the channel conditions, the number of users, and the number of receive antennas at the BS. These findings provide crucial insights when selecting the appropriate value of connected elements in the practical implementation of RDARS-aided systems.
@@
In Fig. 10, we explored the ANMSE performance versus the Rician factor $\kappa_{t}$ under different mode selection schemes with the same phase coefficients, considering both single-user and multiple-user scenarios. In the single-user case, as the Rician factor $\kappa_{t}$ increases from 0 to 1, the channel transitions from Rayleigh fading to a LoS channel. Consequently, the ANMSE performance is gradually improving. Additionally, as the Rician factor grows, the performance gap between the optimized mode selection scheme and the random/fixed mode selection schemes diminishes. This implies that the optimization of the mode selection becomes less critical as the strength of the LoS component increases. On the other hand, the ANMSE performance in the case of multiple users does not exhibit a monotonic relationship with the Rician factor due to the increased interference among multiple users as the LoS components become more dominant. The presence of stronger LoS components will lead to more significant interference, impacting the overall performance. Lastly, when considering only LoS components ($\kappa_{T}=1$), both the optimized mode selection scheme and the random/fixed mode selection scheme achieve the same ANMSE performance. This finding aligns with Proposition 3, as stated in Section III, suggesting the mode selection matrix ${\bf{A}}$ can be arbitrarily chosen when only LoS channels are present. In conclusion, the ANMSE performance in the RDARS is influenced by the Rician factor, with different trends observed in single-user and multiple-user scenarios. The necessity for mode selection optimization diminishes when only LoS components are present. These insights are valuable for guiding the practical implementation of RDARS systems.
V-G ANMSE vs $\beta_{c}$.
Considering such spatially correlated channels, we have studied the ANMSE performance as a function of the number of elements $N$ in the RDARS under various levels of spatial correlation, as illustrated in Fig. 11.
Figure 11: The ANMSE versus the number of elements $N$ in the RDARS under various levels of spatial correlation.
It can be observed from Figs. 11 that an increase in the number of RDARS elements $N$ enhances the MSE performance. However, the spatial correlation at the RDARS negatively affects this performance. As the correlation coefficient $\beta_{c}$ varies from 00 to $1$, the channels associated with the RDARS evolve from uncorrelated Rician channels to fully correlated ones. With an increasing level of spatial correlation, the rank of the channels related to the RDARS decreases. This reduction consequently lowers the spatial multiplexing gain and intensifies multiuser interference, leading to a degradation in performance.
V-H ANSME vs the RDARS’s Position.
Fig. 12 illustrates the ANMSE performance as a function of the horizontal distance of the RDARS, i.e., $d_{R}$, under both single-user and multi-user scenarios. The conventional passive RIS and DAS are compared as benchmarks against the same three-dimensional setup as depicted in Fig. 3. In the single-user system shown in Fig. 12 (a), the passive RIS achieves optimal performance when located near the user or the BS, while the performance of the DAS deteriorates as it approaches the BS. The RDARS demonstrates superior performance when in proximity to the user. However, at a distance of 200 meters, the MSE is lower than at 160 meters when there is one element in connection mode, i.e., $a=1$, due to the dominance of reflection gain over distribution and selection gains. Conversely, when $a=4$, the distribution and selection gains become dominant, reversing this trend. In Fig. 12 (b), the scenario includes four users in the RDARS-aided uplink communication system. The RDARS shows a consistent performance trend with $a=1$ as the horizontal distance varies, driven by the dominant reflection gain. As $a$ increases to 4, the RDARS performance is poorest near the BS. Based on these observations, we conclude that better MSE performance is achieved when the RDARS is closer to the users, offering valuable guidelines for the practical deployment of RDARS-aided systems.
(a) Single user
(b) Multiple user with $M=4$
Figure 12: ANMSE vs the horizontal distance of the RDARS in (a) Single user (b) Multiple users.